### v1

# Differential geometry of particle motion in Stokesian regime

Abstract
We present a differential geometric framework for the motion of a non-Brownian particle in the presence of fixed obstacles in a quiescent fluid, in the deterministic Stokesian regime. While the Helmholtz Minimum Dissipation Theorem suggests that the hydrodynamic resistance tensor $R_{ij}$ acts as the natural Riemannian metric of the fluid domain, we demonstrate that particle trajectories driven by constant external forces are not geodesics of this pure resistance metric. Instead, they experience a geometric drift perpendicular to the geodesic path due to the manifold’s curvature. To reconcile this, we introduce a unified geometric formalism, proving that physical trajectories are geodesics of a conformally scaled metric, $\tilde{g}_{ij}=\mathcal{D}(\mathbf{x})R_{ij}$, where $\mathcal{D}$ is the local power dissipation. This framework establishes that the affine parameter along the trajectory corresponds to the cumulative energy dissipated. We apply this theory to the scattering of a spherical particle by a fixed obstacle, showing that the previously derived trajectory of the particle is recovered as a direct consequence of the curvature of this dissipation-scaled manifold.
preprint: APS/123-QED
The use of geometry to describe physical phenomena is a cornerstone of modern theoretical physics, most notably in Einstein’s description of gravitation [2] and the broader field of geometric mechanics [1, 5]. In these classical frameworks, the physical system is typically conservative, and the logical progression is straightforward: a Lagrangian is defined, the Principle of Least Action yields the equations of motion, and these are interpreted as geodesic equations on a Riemannian manifold.
Here, we extend this geometric treatment to the purely dissipative, Aristotelian domain of low-Reynolds-number hydrodynamics. The Stokesian regime, defined by negligible fluid and particle inertia, is governed by the linear Stokes equations: {align} μ∇^2 u = ∇p,  ∇⋅u = 0. Because the equations are linear and time-reversible, the motion of a suspended rigid particle is governed by a linear mobility relation. For a torque-free particle acted upon by a force $\mathbf{F}$ through its center of buoyancy, the velocity $\mathbf{U}$ is given by:
$\mathbf{F}=\mathbf{R}(\mathbf{x})\cdot\mathbf{U},$ (1)
where $\mathbf{R}(\mathbf{x})$ is the position-dependent hydrodynamic resistance tensor [3].
Historically, the connection between Stokes flow and variational calculus is well-established. Helmholtz [12] proved that of all solenoidal flow fields satisfying given boundary conditions, the Stokes flow minimizes the total power dissipation. Hill and Power [4] extended this to include the rigid-body motion of suspended particles. This minimization principle suggests a compelling geometric hypothesis: Does the hydrodynamic resistance tensor $\mathbf{R}$ act as a Riemannian metric, such that particle trajectories are geodesics of the manifold defined by fluid dissipation?
We demonstrate that while the resistance tensor $\mathbf{R}$ defines the local cost of motion, it is insufficient to describe the motion under constant external forces. We show that the trajectories of sedimenting particles deviate from the geodesics of $\mathbf{R}$ due to a geometric acceleration arising from the manifold’s curvature. We resolve this by introducing a Unified Dissipative Metric, $\tilde{\mathbf{g}}=\mathcal{D}\mathbf{R}$, which scales the resistance geometry by the local power dissipation. We prove that physical trajectories are geodesics of this manifold, establishing a geometric variational principle where the affine parameter corresponds to the cumulative energy dissipated.
I Particle trajectories and Geodesics
We start with the expression for energy dissipated along a trajectory of the particle (see Appendix A for a short derivation):
$E=\int_{0}^{t}\mathbf{R}\cdot\mathbf{U}\cdot\mathbf{U}~d\tau,$ (2)
where the integral is understood to be over a particle trajectory from time $\tau=0$ to $\tau=t$.
Note that the instantaneous particle velocity $\mathbf{U}$ is the tangent vector of the particle trajectory. Therefore, in index notation, equation (2) defines a functional equivalent to the length functional in Riemannian geometry:
$E=\int_{0}^{t}R_{ij}\frac{dx^{i}}{d\tau}\frac{dx^{j}}{d\tau}~d\tau.$
If one applies the Euler-Lagrange equations to extremize this dissipation functional, one obtains the standard homogeneous geodesic equations for the metric $g_{ij}=R_{ij}$. This suggests a compelling hypothesis: Are the physical trajectories of Stokesian particles geodesics of the resistance metric?
To test this, we examine the covariant acceleration of a particle driven by a constant external force $F_{k}$ (e.g., gravity). The equation of motion is given by the mobility relation $U^{i}=g^{ij}F_{j}$. Differentiating this with respect to time yields the kinematic acceleration:
$\frac{dU^{i}}{dt}=\frac{\partial g^{ij}}{\partial x^{k}}U^{k}F_{j}.$ (3)
Using the identity $\partial_{k}g^{ij}=-g^{ia}g^{jb}(\partial_{k}g_{ab})$ and substituting $F_{j}=g_{jl}U^{l}$, we obtain:
$\frac{dU^{i}}{dt}=-g^{ia}(\partial_{k}g_{al})U^{k}U^{l}.$ (4)
The intrinsic geometry of the trajectory, however, is determined by the absolute (covariant) derivative $\frac{DU^{i}}{dt}=\frac{dU^{i}}{dt}+\Gamma^{i}_{jk}U^{j}U^{k}$. Substituting the kinematic acceleration into this definition reveals a non-vanishing term:
$\frac{DU^{i}}{dt}=-\frac{1}{2}g^{ia}\left(\partial_{a}g_{jk}\right)U^{j}U^{k}\neq 0.$ (5)
This result contradicts the hypothesis that the trajectory is a geodesic of the resistance metric alone. While the integral (2) describes the cost of motion, the physical particle driven by a constant Euclidean force does not traverse the path of minimum resistance. Instead, it experiences a geometric “drift” perpendicular to the geodesic path, necessitating a more unified geometric treatment.
II The Jacobi-Maupertuis Resolution
The deviation observed in Equation (5) arises because the resistance metric $g_{ij}=R_{ij}$ describes the geometry of the medium, defining the instantaneous cost of motion, but ignores the dynamics of the driving force. To reconcile the kinematics with the geometry, we invoke a dissipative analogue to the Jacobi-Maupertuis principle.
In classical mechanics, the Jacobi-Maupertuis metric $\tilde{g}_{ij}=2(E-V)g_{ij}$ transforms a dynamical problem into a geometric one by weighing the kinetic energy metric with the available potential energy. For a Stokesian particle, the governing variational principle is the minimization of total energy dissipation [12, 7]. The particle trajectory is the path that minimizes the action functional:
$S=\int_{0}^{T}\mathcal{D}(\mathbf{x},\dot{\mathbf{x}})\,dt,$ (6)
where $\mathcal{D}=\mathbf{U}\cdot\mathbf{R}\cdot\mathbf{U}=R_{ij}\dot{x}^{i}\dot{x}^{j}$ is the instantaneous power dissipation.
To interpret this physical minimization as a geometric geodesic problem, we seek a Riemannian manifold where the action $S$ corresponds to the arc length. We introduce a Unified Dissipative Metric $\tilde{\mathbf{g}}$ defined by a conformal scaling of the resistance tensor:
$\tilde{g}_{ij}(\mathbf{x})=\mathcal{D}(\mathbf{x})R_{ij}(\mathbf{x}).$ (7)
Using this metric, the squared norm of the velocity vector becomes $\tilde{g}_{ij}\dot{x}^{i}\dot{x}^{j}=\mathcal{D}(R_{ij}\dot{x}^{i}\dot{x}^{j})=\mathcal{D}^{2}$. Consequently, the local dissipation rate can be expressed purely geometrically as $\mathcal{D}=\sqrt{\tilde{g}_{ij}\dot{x}^{i}\dot{x}^{j}}$. Substituting this into the action functional yields the standard Riemannian length functional:
$S=\int_{0}^{T}\sqrt{\tilde{g}_{ij}\frac{dx^{i}}{dt}\frac{dx^{j}}{dt}}\,dt.$ (8)
By definition, the path that minimizes this functional is a geodesic of the metric $\tilde{\mathbf{g}}$.
This resolves the discrepancy found in the pure resistance formulation. While the resistance tensor $R_{ij}$ creates a geometric “drift”, the scaling factor $\mathcal{D}(\mathbf{x})$ introduces a counter-acting “conformal force.” As derived explicitly in Appendix D, the gradient of the dissipation rate exactly cancels the geometric drift arising from the resistance tensor.
The resulting geodesic equation is homogeneous only when parameterized by the natural arc length $s$, which satisfies $ds=\mathcal{D}dt$. This offers a profound physical interpretation: the affine parameter measuring “distance” along the trajectory is the cumulative energy dissipated. Thus, a Stokesian particle under a constant force traces a geodesic in the dissipation-scaled manifold, effectively following the path of stationary dissipative action.
III Application to a two-sphere system
Risbud and Drazer [9] have derived the equations describing the trajectory of a sphere moving past another fixed sphere in an unbounded fluid, in the Stokesian regime:
$b_{in}=y\exp{H(r)},$ (9)
where $b_{in}$ is a constant (the impact parameter), $y$ is the ordinate, and $r$ is the radial coordinate. The function $H(r)$ is given by:
$H(r)=\displaystyle\int_{r}^{\infty}\frac{R_{A}(s)-R_{B}(s)}{sR_{A}(s)}ds.$
Given the symmetries of the problem, the functions $R_{A}(r)$ and $R_{B}(r)$ constitute the hydrodynamic resistance tensor [9]:
[…]
where $\mathbf{\hat{r}}$ is the unit vector along the line joining the two centers pointing towards the moving sphere.
From the formalism established in the previous section, we identify equation (9) as the geodesic equation on the manifold defined by the Unified Dissipative Metric $\tilde{\mathbf{g}}=\mathcal{D}(\mathbf{x})\mathbf{R}(\mathbf{x})$. The analytical solution represents the path that minimizes the dissipative arc length.
The function $H(r)$ serves as a measure of the anisotropic curvature of this manifold. To see this physically, consider the isotropic limit where $R_{A}(r)=R_{B}(r)=R(r)$. In this case, the mobility is isotropic, meaning the velocity is always parallel to the driving force ($\mathbf{U}\parallel\mathbf{F}$). The local dissipation rate becomes $\mathcal{D}=F^{2}/R(r)$. Consequently, the metric simplifies to:
$\tilde{\mathbf{g}}=\left(\frac{F^{2}}{R}\right)R\bm{\delta}=F^{2}\bm{\delta}.$ (11)
Since the metric is proportional to the identity tensor (conformal to Euclidean space with a constant factor), the Christoffel symbols vanish, and the geodesics are straight lines. This matches the analytical result: if $R_{A}=R_{B}$, the integral vanishes, $H(r)=0$, and the trajectory becomes $y=b_{in}$ (a straight line).
However, in the presence of hydrodynamic interactions, $R_{A}\neq R_{B}$. The factor $(R_{A}-R_{B})/R_{A}$ acts as a geometric “refractive index” gradient. The exponential term $\exp{H(r)}$ quantifies how the particle is effectively steered by the curvature of the dissipation manifold.
Crucially, this result highlights the necessity of the dissipation scaling. Had we calculated the geodesic of the pure resistance metric $\mathbf{R}$ (as one might for a particle dragged by optical tweezers), the trajectory would differ. The pure resistance geodesic minimizes the accumulation of $R_{ij}$, whereas the physical trajectory minimizes the accumulation of $\mathcal{D}$. The factor $\mathcal{D}$ acts as the necessary counter-weight to the resistance gradient, ensuring that the particle follows the path prescribed by the constant external force.
IV Discussion and Conclusions
We have presented a differential geometric framework that resolves the kinematics of Stokesian particles driven by constant external forces. Our central result is the identification of the Unified Dissipative Metric, $\tilde{\mathbf{g}}=\mathcal{D}(\mathbf{x})\mathbf{R}(\mathbf{x})$, as the natural geometry of the motion. While the hydrodynamic resistance tensor $\mathbf{R}$ defines the local cost of motion, it is the dissipation-scaled metric $\tilde{\mathbf{g}}$ that governs the global trajectory. Previously derived trajectory of a sphere moving past an obstacle [9] is thus revealed to be a geodesic in this dissipation-manifold, where the affine parameter corresponds to the cumulative energy dissipated.
IV.1 Relationship to Geometric Swimming
This framework parallels, yet remains distinct from, the gauge-theoretic description of self-propulsion established by Shapere and Wilczek [11, 10] and the general relativistic treatment of ‘swimming in spacetime’ by Wisdom [13]. Shapere and Wilczek encode the hydrodynamics in a gauge potential over the manifold of shapes, where the optimal swimming stroke corresponds to a geodesic in shape space. Wisdom refers to this work and follows a similar line of reasoning in the context of locomotion in spacetime.
The fundamental distinction lies in the source of motion. Both works address active matter, where internal degrees of freedom (cyclic shape deformations) couple to the fluid to generate thrust. In contrast, our framework addresses passive matter driven by external fields (gravity, electrophoresis) in the physical embedding space. Despite this difference, both frameworks share a common physical origin: the minimization of viscous dissipation. The “efficiency metric” on shape space is mathematically dual to the “resistance metric” on physical space. Our unified metric $\tilde{\mathbf{g}}$ can thus be viewed as the passive-transport counterpart to the active-swimming gauge potential, completing the geometric picture of Stokesian dynamics.
IV.2 The Duality of Stokesian Metrics
Our results suggest a useful classification for experimentalists involving two distinct classes of optimal paths:
  1. 1.
    Resistance Geodesics ($\mathbf{R}$): These are the paths of Global Minimum Dissipation for a particle transported between two points. These paths can be realized using active control methods, such as optical tweezers, where the controller minimizes the total work done against the fluid.
  2. 2.
    Dissipative Geodesics ($\tilde{\mathbf{g}}$): These are the Natural Trajectories of particles sedimenting under constant forces. Here, the particle balances the local resistance gradient against the driving power, following a path of stationary dissipative action.
The geometric drift derived in this Letter — the “extraneous term” $\frac{1}{2}g^{ia}\left(\partial_{a}g_{jk}\right)U^{j}U^{k}$ from Equation 5— is precisely the penalty paid by the natural trajectory for deviating from the resistance geodesic due to the constraint of a fixed external force direction.
IV.3 Broader Implications
The concept of the Unified Dissipative Metric extends beyond simple sedimentation. Any low-Reynolds-number system where the local dissipation rate $\mathcal{D}$ can be defined as a function of position—including electrophoresis, magnetophoresis, or motion through porous media—can be mapped onto a Riemannian manifold. This opens the door to “Geometric Microfluidics,” where obstacle geometries are designed specifically to curve the effective spacetime of the channel, focusing or sorting particles based on their anisotropic resistance signatures.
Acknowledgements.
The author is grateful to Prof. German Drazer of Rutgers University, Prof. Ravindra Kulkarni of Bhaskaracharya Pratishthana (Pune) , and Prof. Rahul Kashyap of IIT Bombay for enlightening discussions.
Appendix A Power dissipated by a single moving particle in a quiescent fluid
Consider a single torque-free particle moving under the action of a constant force $\mathbf{F}$ acting on its center of buoyancy in a quiescent unbounded fluid, in the presence of fixed boundaries. The viscous dissipation in this case would be:
$\dot{E}=\int_{V_{p}}\mathbf{\sigma}\cdot\mathbf{e}~dV,$ (12)
where, $V_{p}$ is the volume of the particle, $\mathbf{\sigma}$ is the stress tensor, and $\mathbf{e}=\frac{1}{2}\left(\nabla\mathbf{u}+\left(\nabla\mathbf{u}\right)^{T}\right)$. Using Stokes theorem to convert the volume integral to a surface integral over the particle surface, we obtain
$P=\dot{E}=\mathbf{F}\cdot\mathbf{U}=\mathbf{R}\cdot\mathbf{U}\cdot\mathbf{U}.$ (13)
In equation (13), only translational velocity and force contribute to the dissipated power, because the particle is known to be torque-free. Also, the information about the geometry of the problem is captured by the hydrodynamic resistance tensor $\mathbf{R}$ [3, 8].
Consequently, the energy dissipated along a trajectory of the particle is
$E=\int_{0}^{t}\mathbf{R}\cdot\mathbf{U}\cdot\mathbf{U}~d\tau,$ (14)
where, the integral is understood to be over a particle trajectory from time $\tau=0$ to $\tau=t$.
Appendix B Geodesics on a manifold
A geodesic on a manifold is a curve that locally minimizes ‘distance’ on the manifold, as defined by a metric tensor. Formally, between points $P_{1}$ and $P_{2}$ on a smooth manifold $\mathcal{M}$ a geodesic curve minimizes
$\ell_{P_{1},P_{2}}=\int_{s_{1}}^{s_{2}}~ds,$ (15)
where, dsds is the line element along the curve, defined as the quadratic form
$ds^{2}=g_{ij}dx^{i}dx^{j},$ (16)
if $g_{ij}$ is the metric tensor on a manifold $\mathcal{M}$.
Here, we are using the Einstein summation notation such that summation is assumed over repeated indices. Further, indices in subscript signify covariance while those in superscript signify contravariance [6]. The local coordinates are $x^{j}$, where $j$ varies from 0 to the number of dimensions of the space in which $\mathcal{M}$ is embedded.
For a particular scalar parameter $\alpha$ that parametrizes the curve, the quadratic form (16) can be rewritten as
$L(\alpha)^{2}:=\left(\frac{ds(\alpha)}{d\alpha}\right)^{2}=g_{ij}\frac{dx^{i}(\alpha)}{d\alpha}\frac{dx^{j}(\alpha)}{d\alpha}.$ (17)
Therefore, in terms of a particular parametrization $\alpha$, the length between points $P_{1}$ and $P_{2}$ in equation (15) is {align} ℓ_P_1, P_2 = ∫_α_1^α_2 dsdα dα
= ∫_α_1^α_2 L(α)  dα
= ∫_α_1^α_2 g_ijdxi(α)dαdxj(α)dα dα.
Writing the Euler-Lagrange equations corresponding to equation (B) for Lagrangian $L(\alpha)$ give us the following (inhomogeneous) geodesic equation (Ch. 17, eq. 17.7 in [6]): {align} d2xidα2 + g^ij∂gjk∂xm dxkdα dxmdα - 12g^ij∂gkm∂xj dxkdα dxmdα
= 12dxidα(∂gkm∂xpdxkdα dxmdα dxpdα).
Instead, we change the parametrization to some other scalar parameter $\beta$, such that we can define our Lagrangian as,
$\tilde{L}\left(\beta\right):=g_{ij}\frac{dx^{i}(\beta)}{d\beta}\frac{dx^{j}(\beta)}{d\beta},$ (18)
and consider the integral,
$\int_{\beta_{1}}^{\beta_{2}}\tilde{L}(\beta)~d\beta=\int_{\beta_{1}}^{\beta_{2}}g_{ij}\frac{dx^{i}(\beta)}{d\beta}\frac{dx^{j}(\beta)}{d\beta}~d\beta.$ (19)
Note that the same curves on $\mathcal{M}$ minimize (19) as those which minimize (B), just with a different parametrization.
The Euler-Lagrange equations for (19) give us the homogeneous version of the geodesic equation,
$\frac{d^{2}x^{i}}{d\beta^{2}}+g^{ij}\frac{\partial g_{jk}}{\partial x^{m}}\frac{dx^{k}}{d\beta}\frac{dx^{m}}{d\beta}-\frac{1}{2}g^{ij}\frac{\partial g_{km}}{\partial x^{j}}\frac{dx^{k}}{d\beta}\frac{dx^{m}}{d\beta}=0,$ (20)
and the parameter $\beta$ is called an affine parameter (Ch. 9, Exercise 9.2 in [6]).
The parameters $\alpha$ and $\beta$ are related to each other as:
$\frac{d\beta}{d\alpha}=L(\alpha).$ (21)
Therefore, if we know that an integral of the form (19) is minimized along a curve on a manifold, then we can simply read-out the corresponding metric tensor and the affine parameter along the curve.
Appendix C Proof of Geometric Consistency
In this section, we explicitly prove the claim that the physical trajectory of a Stokesian particle driven by a potential force field $\mathbf{F}(\mathbf{x})=-\nabla\Phi$ corresponds to a geodesic of the Jacobi-Maupertuis metric $\tilde{g}_{ij}=\mathcal{D}\left(\mathbf{x}\right)R_{ij}$, with $\mathcal{D}$ being the dissipation.
C.1 Transformation of the Covariant Acceleration
Consider a conformally scaled metric $\tilde{g}_{ij}=\Omega(\mathbf{x})R_{ij}$. We investigate the covariant acceleration of the particle trajectory with respect to this metric. The Christoffel symbols transform as:
$\tilde{\Gamma}^{i}_{jk}=\Gamma^{i}_{jk}+\frac{1}{2\Omega}\left(\delta^{i}_{k}\partial_{j}\Omega+\delta^{i}_{j}\partial_{k}\Omega-g_{jk}g^{im}\partial_{m}\Omega\right).$ (22)
The total covariant acceleration is:
$\frac{\tilde{D}U^{i}}{dt}=\frac{dU^{i}}{dt}+\tilde{\Gamma}^{i}_{jk}U^{j}U^{k}.$ (23)
Substituting the split connection $\tilde{\Gamma}=\Gamma+C$: {align} ~DUidt = ⏟( dUidt+ Γ^i_jk U^j U^k )_Resistance Accel. DU^i/dt
+ 12Ω ( 2 U^i (U^k ∂_k Ω) - (g_jk U^j U^k) g^im ∂_m Ω). We identify the scalar term $U^{k}\partial_{k}\Omega=\dot{\Omega}$ (rate of change along path) and the dissipation term $g_{jk}U^{j}U^{k}=\mathcal{D}$.
$\frac{\tilde{D}U^{i}}{dt}=\frac{DU^{i}}{dt}+\frac{\dot{\Omega}}{\Omega}U^{i}-\frac{\mathcal{D}}{2\Omega}g^{im}\partial_{m}\Omega.$ (24)
C.2 The Dissipation Metric Proof
We previously established that the acceleration in the pure resistance metric is driven by the gradient of dissipation. Specifically, for constant force motion where $\partial_{m}\mathcal{D}=-U^{j}U^{k}\partial_{m}g_{jk}$, we found:
$\frac{DU^{i}}{dt}=-\frac{1}{2}g^{im}(\partial_{m}g_{jk})U^{j}U^{k}=+\frac{1}{2}g^{im}\partial_{m}\mathcal{D}.$ (25)
Substituting this result into Eq. (24):
$\frac{\tilde{D}U^{i}}{dt}=\frac{1}{2}g^{im}\partial_{m}\mathcal{D}+\frac{\dot{\Omega}}{\Omega}U^{i}-\frac{\mathcal{D}}{2\Omega}g^{im}\partial_{m}\Omega.$ (26)
Grouping the gradient terms:
$\frac{\tilde{D}U^{i}}{dt}=\frac{1}{2}g^{im}\left(\partial_{m}\mathcal{D}-\frac{\mathcal{D}}{\Omega}\partial_{m}\Omega\right)+\frac{\dot{\Omega}}{\Omega}U^{i}.$ (27)
For the trajectory to be a geodesic, the non-tangential (transverse) acceleration must vanish. This requires the term in the parentheses to be zero:
[…]
Integrating this condition yields the required scaling factor:
[…]
Thus, by choosing the metric scaling $\Omega=\mathcal{D}$, the transverse drift is exactly cancelled. The equation of motion reduces to:
$\frac{\tilde{D}U^{i}}{dt}=\frac{\dot{\mathcal{D}}}{\mathcal{D}}U^{i}.$ (30)
This describes a projective geodesic: the particle follows the geometric path of a geodesic, with a parameterization speed determined by the local dissipation rate. The trajectory shape depends only on the metric $\tilde{\mathbf{g}}=\mathcal{D}\mathbf{R}$.
C.3 Affine Parameterization and Energy Dissipation
In the previous step, we derived the equation of motion in the metric $\tilde{g}_{ij}=\mathcal{D}R_{ij}$:
$\frac{\tilde{D}U^{i}}{dt}=\frac{\dot{\mathcal{D}}}{\mathcal{D}}U^{i}.$ (31)
The presence of the term proportional to $U^{i}$ indicates that while the trajectory traces a geodesic curve, the physical time $t$ is not an affine parameter. To transform this into the standard geodesic equation $\frac{\tilde{D}V^{i}}{ds}=0$, we introduce a new parameter $s(t)$ such that the tangent vector is $V^{i}=dx^{i}/ds$.
Using the relation $U^{i}=\dot{s}V^{i}$ (where $\dot{s}=ds/dt$), the covariant acceleration expands as:
$\frac{\tilde{D}U^{i}}{dt}=\frac{\tilde{D}(\dot{s}V^{i})}{dt}=\ddot{s}V^{i}+\dot{s}^{2}\frac{\tilde{D}V^{i}}{ds}.$ (32)
Equating this to the right-hand side of the pre-geodesic equation ($\frac{\dot{\mathcal{D}}}{\mathcal{D}}\dot{s}V^{i}$) and enforcing the geodesic condition $\frac{\tilde{D}V^{i}}{ds}=0$, we obtain the scalar differential equation for the parameter $s$:
[…]
Integrating yields the relation between the parameter and physical time:
[…]
Since $\mathcal{D}$ represents the rate of energy dissipation, the differential dsds corresponds to the incremental energy dissipated along the path.
Conclusion: The trajectory of a Stokesian particle under a constant force is a geodesic of the metric $\tilde{\mathbf{g}}=\mathcal{D}(\mathbf{x})\mathbf{R}(\mathbf{x})$, parameterized by the cumulative dissipated energy.
C.4 Variational Derivation of the Geodesic Equation
We explicitly show that minimizing the total energy dissipation is equivalent to minimizing the arc length on the manifold $\tilde{\mathcal{M}}$ with metric $\tilde{g}_{ij}=\mathcal{D}R_{ij}$.
The total energy dissipated along a trajectory is given by the functional:
$S=\int_{0}^{T}\mathcal{D}(\mathbf{x},\dot{\mathbf{x}})\,dt=\int_{0}^{T}\left(R_{ij}\frac{dx^{i}}{dt}\frac{dx^{j}}{dt}\right)dt.$ (35)
We introduce the metric $\tilde{g}_{ij}=\mathcal{D}R_{ij}$. We verify the identity relating the dissipation rate to the line element of this metric:
$\tilde{g}_{ij}\frac{dx^{i}}{dt}\frac{dx^{j}}{dt}=(\mathcal{D}R_{ij})U^{i}U^{j}=\mathcal{D}(R_{ij}U^{i}U^{j})=\mathcal{D}^{2}.$ (36)
Thus, the instantaneous dissipation rate can be written as:
$\mathcal{D}=\sqrt{\tilde{g}_{ij}\dot{x}^{i}\dot{x}^{j}}.$ (37)
Substituting this back into the action functional:
$S=\int_{0}^{T}\sqrt{\tilde{g}_{ij}\frac{dx^{i}}{dt}\frac{dx^{j}}{dt}}\,dt.$ (38)
This is the standard definition of the path length functional in Riemannian geometry. By the fundamental theorem of variational calculus, the extrema of this functional are geodesics of the metric $\tilde{g}_{ij}$.
To recover the homogeneous equation of motion, we define the affine parameter $s$ via the arc length:
$ds=\sqrt{\tilde{g}_{ij}dx^{i}dx^{j}}=\mathcal{D}\,dt.$ (39)
In terms of $s$, the Euler-Lagrange equations for the functional $S=\int ds$ yield the standard homogeneous geodesic equation:
$\frac{\tilde{D}}{ds}\left(\frac{dx^{k}}{ds}\right)=\frac{d^{2}x^{k}}{ds^{2}}+\tilde{\Gamma}^{k}_{ij}\frac{dx^{i}}{ds}\frac{dx^{j}}{ds}=0.$ (40)
This confirms that the physical requirement of minimum dissipation necessitates that the particle follows a geodesic in the dissipation-scaled manifold.
Appendix D Derivation of the Two-Sphere Solution
In this appendix, we explicitly re-derive the Risbud-Drazer trajectory, $b_{in}=y\exp{H(r)}$. Given that it is a system under Stokesian regime and the particle is moving under the action of a constant force, it trajectory is already proven to be a geodesic on the corresponding dissipation manifold.
It remains to solve the mobility equation for a particle driven by a force $\mathbf{F}$:
$\mathbf{U}=\mathbf{R}^{-1}\cdot\mathbf{F}.$ (41)
We define the geometry using polar coordinates $(r,\theta)$ centered on the fixed sphere. The resistance tensor $\mathbf{R}$ is diagonal in this basis:
[…]
The external force $\mathbf{F}=F\mathbf{\hat{x}}$ is aligned with the symmetry axis. Its components in polar coordinates are:
$F_{r}=F\cos\theta,\quad F_{\theta}=-F\sin\theta.$ (43)
Using the mobility relation (inverse resistance), the velocity components are:
$U_{r}=\frac{F_{r}}{R_{A}(r)}=\frac{F\cos\theta}{R_{A}(r)},\quad U_{\theta}=\frac{F_{\theta}}{R_{B}(r)}=\frac{-F\sin\theta}{R_{B}(r)}.$ (44)
The differential equation of the trajectory is given by the ratio of the velocity components:
$\frac{1}{r}\frac{dr}{d\theta}=\frac{U_{r}}{U_{\theta}}=-\frac{R_{B}(r)}{R_{A}(r)}\cot\theta.$ (45)
Rearranging to separate variables yields:
$\frac{R_{A}(r)}{R_{B}(r)}\frac{dr}{r}=-\cot\theta\,d\theta=-d(\ln\sin\theta).$ (46)
To recover the form of the analytical solution, we transform coordinates from $(r,\theta)$ to $(r,y)$, utilizing the relation $y=r\sin\theta$. Taking the differential of the logarithm:
[…]
Substituting this geometric identity back into the separated differential equation:
$\frac{R_{A}(r)}{R_{B}(r)}\frac{dr}{r}=\frac{dr}{r}-\frac{dy}{y}.$ (48)
Rearranging to isolate the vertical coordinate $y$:
$\frac{dy}{y}=\left(1-\frac{R_{A}(r)}{R_{B}(r)}\right)\frac{dr}{r}=\frac{R_{B}(r)-R_{A}(r)}{R_{B}(r)}\frac{dr}{r}.$ (49)
We integrate this expression from the current position $r$ to infinity. At the far-field limit ($r\to\infty$), the vertical position $y$ approaches the impact parameter $b_{in}$.
$\int_{y}^{b_{in}}d(\ln y^{\prime})=\int_{r}^{\infty}\frac{R_{B}(s)-R_{A}(s)}{sR_{B}(s)}ds.$ (50)
The left-hand side evaluates to $\ln(b_{in}/y)$. By defining the anisotropic curvature function $H(r)$ consistently with the mobility formulation:
$H(r)\equiv\int_{r}^{\infty}\frac{R_{B}(s)-R_{A}(s)}{sR_{B}(s)}ds,$ (51)
we obtain the final trajectory equation:
[…]
which is the equation of a trajectory we sought to derive.